# Cryogenic cesium iodide as a potential PET material

### v2

@@ 1 Introduction
Therefore, developing a cost-effective TBPET system will require identifying a less expensive alternative to lutetium-based crystals, ideally one that provides similar performance. In this context, the possibility of using scintillating crystals based on cesium iodide (CsI) is worth being explored. Table 1 shows the properties of CsI compared with other crystals commonly used for PET.
Material | (cm) | (g/cm3) | LY (/keV) | (ns) | Peak (nm) | Cost (€/cm3) | |
LYSO | 66 | 1.14 | 7.4 | 30 | 53 | 420 | 45 |
BGO | 74 | 1.12 | 7.1 | 10 | 300 | 480 | 25 |
CsI(Tl) | 54 | 1.86 | 4.5 | 55 | 1000 | 560 | 5 |
CsI (300 K) | 54 | 1.86 | 4.5 | 5 | 15 | 310 | 5 |
CsI (100 K) | 54 | 1.86 | 4.5 | 100 | 800 | 350 | 5 |
@@ 2 Experimental setup
We set up a table-top experiment, illustrated in figure 1, employing a pair of pure CsI crystals of dimensions mm3 provided by AMCRYS and polished on all sides. The crystals were wrapped with 3 layers of 80 -thick polytetrafluoroethylene (PTFE) tape and mounted on a copper plate which can be put in contact with a liquid nitrogen dewar through a cold finger. A Na-22 radioactive source with an activity of 40 kBq was placed between the two crystals. The light was detected by a pair of mm2 Hamamatsu S13360-3025CS MPPCs [12]. An aluminum infrared shield was placed above the copper plate to minimize radiative losses. The signals were acquired by a Tektronix MSO44 oscilloscope triggering the coincidence of the photosensors with a sampling rate of 3.125 Gs/s. Full waveforms were saved to disk with a trigger rate of approximately 5 Hz.
@@ 3.1 Light yield and energy resolution
where is the number of photoelectrons, is the photon detection efficiency of the MPPC, and is the detection efficiency of the setup. The weighted average PDE for the room temperature CsI emission spectrum is 12%, compared with the 17% at cryogenic temperature (see figure 2).
@@
The voltage breakdown of MPPCs is proportional to the temperature with a slope of approximately 54 mV/K [12]. At K, we adjusted the bias voltage by reducing it by 10.3 V to compensate for this effect. We then repeated the single photon measurement. The integration window was increased to 5.5 , due to the longer decay time of CsI at low temperatures (see section 3.2). The MPPC dark count rate is halved every 5.3 K [16], becoming negligible for our integration time at K. Under these low-temperature conditions, the number of photoelectrons increased by over an order of magnitude. This corresponds to light yields of photons/MeV and photons/MeV for the respective crystals.
@@ 4 Summary and future prospects
We have characterized the response of a pair of pure CsI crystals read out by solid state photosensors (Hamamatsu MPPCs) in a setup cooled at cryogenic temperatures. At K, an energy resolution of 6.3% and a coincidence time resolution of 1.84 ns have been achieved. The light yield and decay time have been measured as a function of temperature, with results in good agreement with the existing literature.
The energy resolution of cryogenic CsI is significantly better than the one of the scintillators used in modern PET scanner (LYSO, BGO), which is typically above 10% [2]. The coincidence time resolution, while sufficient for PET applications, is too large to support time-of-flight (TOF) measurements. Thus, this material might be a promising candidate for TBPET scanners, where the scintillator cost represents a large fraction of the total cost and can potentially offset the added complexity of a cryogenic system.
A full simulation of a PET scanner based on cryogenic CsI crystals will be developed, in order to evaluate the effect of an improved energy resolution and a lack of TOF on the image quality. Experimentally, the setup will be scaled up to two matrices of 64 SiPMs each, in order to test the electronics and the cryogenic system needed for a larger number of channels.

### v3

@@ 1 Introduction
Therefore, developing a cost-effective TBPET system will require identifying a less expensive alternative to lutetium-based crystals, ideally one that provides similar performance. In this context, the possibility of using scintillating crystals based on cesium iodide (CsI) is worth being explored. Table 1 shows the properties of CsI compared with other crystals commonly used for PET. At parity of radiation length, CsI is five times less expensive than LYSO and two-three times less expensive than BGO. Additionally, its raw materials are widely available, and crystal growth requires a significantly lower temperature – approximately 800 ∘C, compared to the 1000 ∘C and 2000 ∘C needed for BGO and LYSO, respectively [4].
| Material | (cm) | (g/cm3) | LY (/keV) | (ns) | Peak (nm) | Cost (€/cm) | |
| LYSO | 66 | 1.14 | 7.4 | 30 | 53 | 420 | 50 |
| BGO | 74 | 1.12 | 7.1 | 10 | 300 | 480 | 28 |
| CsI(Tl) | 54 | 1.86 | 4.5 | 55 | 1000 | 560 | 10 |
| CsI (300 K) | 54 | 1.86 | 4.5 | 5 | 15 | 310 | 10 |
| CsI (100 K) | 54 | 1.86 | 4.5 | 100 | 800 | 350 | 10 |
@@ 2 Experimental setup
We set up a table-top experiment, illustrated in figure 1, employing a pair of pure CsI crystals of dimensions mm3 provided by AMCRYS and polished on all sides. Their length corresponds to approximately 1.1 . The crystals were wrapped with 3 layers of 80 -thick polytetrafluoroethylene (PTFE) tape and mounted on a copper plate which can be put in contact with a liquid nitrogen dewar through a cold finger. A Na-22 radioactive source with an activity of 40 kBq was placed between the two crystals. The light was detected by a pair of mm2 Hamamatsu S13360-3025CS MPPCs [13]. An aluminum infrared shield was placed above the copper plate to minimize radiative losses. The signals were acquired by a Tektronix MSO44 oscilloscope triggering the coincidence of the photosensors with a sampling rate of 3.125 Gs/s and a bandwidth of 200 MHz. Full waveforms were saved to disk with a trigger rate of approximately 5 Hz.
@@ 3.1 Light yield and energy resolution
where is the number of photoelectrons, is the photon detection efficiency of the MPPC, and is the detection efficiency of the setup. The weighted average PDE for the room temperature CsI emission spectrum is 12%, compared with the 17% for the CsI cryogenic spectrum (see figure 2). However, conflicting measurements exist regarding the PDE of Hamamatsu MPPCs at cryogenic temperatures [14, 15], and no measurements is available for our specific model. Thus, a dedicated study is needed to improve the precision of this parameter for the estimation of the absolute light yield.
@@
The voltage breakdown of MPPCs is proportional to the temperature with a slope of approximately 54 mV/K [13]. At K, we adjusted the bias voltage by reducing it by 10.3 V to compensate for this effect. We then repeated the single photon measurement. The integration window was increased to 5.5 , due to the longer decay time of CsI at low temperatures (see section 3.2). The MPPC dark count rate is halved every 5.3 K [19], becoming negligible for our integration time at K. Under these low-temperature conditions, the number of photoelectrons increased by over an order of magnitude. This corresponds to light yields of photons/MeV and photons/MeV for the respective crystals. The response of the photosensor was verified to be linear by triggering only one channel and measuring the amount of charge at the 1275 keV peak.
@@ 4 Summary and future prospects
We have characterized the response of a pair of pure CsI crystals read out by solid state photosensors (Hamamatsu MPPCs) in a setup cooled at cryogenic temperatures. At K, an energy resolution of 6.3% and a coincidence time resolution of 1.84 ns have been achieved. The light yield and decay time have been measured as a function of temperature, with results in good agreement with the existing literature. For comparison, LYSO typically achieves energy resolutions around 7–10% [24], while BGO-based systems, such as those from GE [25], can reach approximately 9% at the system level.
The coincidence time resolution, while sufficient for PET applications, is too large to support time-of-flight (TOF) measurements. Thus, this material might be a promising candidate for TBPET scanners, where the scintillator cost represents a large fraction of the total cost and can potentially offset the added complexity of a cryogenic system.
A full simulation of a PET scanner based on cryogenic CsI crystals will be developed, in order to evaluate the effect of an improved energy resolution and a lack of TOF on the image quality. Experimentally, the setup will be scaled up to two matrices of 64 SiPMs each, in order to test the electronics and the cryogenic system needed for a larger number of channels. Monolithic crystals will also be tested. This represents an interesting option, as it could enable the measurement of depth-of-interaction by using the spatial distribution of the collected light, thereby mitigating the drawback of the larger of the CsI.
Beyond its potential in PET, the properties of cryogenic CsI suggest its applicability to other medical imaging modalities. For instance, Single Photon Emission Computed Tomography (SPECT) could benefit from the enhanced spectral performance of cryogenic CsI. While the prolonged pulse duration increases detector dead time, potentially limiting count rates, this trade-off might be acceptable in SPECT systems where high energy discrimination is prioritized over timing. Further studies are needed to evaluate its practical implementation, including the optimization of readout electronics to mitigate the impact of the long decay time.